# Flag complex face structures and decompositions

### v1

@@ Abstract
We consider how $f$-vectors of Cohen–Macaulay and vertex decomposable flag complexes in geometric settings decompose and interact with geometric transformations. This includes subdivisions of simplicial complexes and explicit equations connecting them to $f$-vectors. Applications from this approach include positivity properties of polynomials built out of simplicial complexes lying between unimodality and real-rootedness.
@@ Introduction
While the $h$-vectors of simplicial complexes have many well-known algebraic and geometric structures, their $f$-vectors often do not readily admit structural properties that apply in general. We use the case of flag simplicial complexes satisfying natural algebraic properties to study how they decompose and interact with geometric transformations. Along the way, we will see connections to positivity questions on reciprocal/palindromic polynomials.
The first structural property we will consider is Boolean decompositions of simplicial complexes (Definition 1.2), which come up in the context of the positivity questions mentioned above. The specific motivating result is the following:
Proposition 0.1.
(Nevo–Petersen–Tenner, Proposition 6.2 on p. 1377 of [24])
Note that the part involving $|F|$ is multiplied by $2$.
The positivity question mentioned above is gamma positivity surveyed in [3], which connects a wide variety of topics such as permutation statistics, triangulations of simplicial complexes, and the Hopf conjecture for certain piecewise Euclidean manifolds.
In this direction, we obtain some results on how these decompositions interact with transformations and recursive properties of flag simplicial complexes.
Theorem 0.2.
  1. 1.
    (Proposition 1.5) The edge subdivision of a simplicial complex that has a Boolean decomposition (Definition 1.2) has a generalized Boolean decomposition (Definition 1.4).
  2. 2.
    (Part 1 of Corollary 1.7) Suppose that a Cohen–Macaulay simplicial complex $\Delta$ has an associated simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition. Then, the same is true for an edge subdivision $\Delta^{\prime}$ of $\Delta$. In other words, there is a simplicial complex $\Gamma^{\prime}$ such that $h(\Delta^{\prime})=f(\Gamma^{\prime})$ and $\Gamma^{\prime}$ has a Boolean decomposition.
  3. 3.
    • $\Delta$ is vertex decomposable and flag.
    • $\Delta$ is a balanced, vertex decomposable, flag simplicial complex.
    Any simplicial complexes associated to antistars and links of vertices with Boolean decompositions can be used to construct a flag (resp. quasi-flag from Definition 4.2 on p. 99 of [10]) simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition.
Next, we consider connections between structures of $f$-vectors and positivity properties of reciprocal/palindromic polynomials. The main tool we use is an interpretation of the gamma polynomial in terms of an inverted Chebyshev expansion which we found in [26]. This was motivated by thinking about gamma polynomial-like invariants of the “usual” basis of reciprocal/palindromic polynomials of a given degree.
@@
This enables us to apply the following result of Hetyei on compatibility between a variant of the usual $f$-polynomial and Tchebyshev transforms $T:\mathbb{R}[x]\longrightarrow\mathbb{R}[x]$ sending $x^{n}\mapsto T_{n}(x)$.
@@
As a result, we find a connection with signed unused color complexes (Definition 2.19) of balanced simplicial complexes (Theorem 2.22).
@@
Let $\Delta$ be a flag simplicial homology sphere of even dimension $d$ and
$[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{k-1}(\mathcal{D}(\Gamma))$
and
Substituting in $\alpha=\frac{\beta-1}{2}$, we have that
where
$F(u)\coloneq\sum_{i=0}^{D}f_{i-1}(A)\left(\frac{u-1}{2}\right)^{i}$
for a cell complex $A$ of dimension $D-1$.
Combining this with Theorem 0.4, we find that the the $f$-polynomial of the interval poset of $S$ is a scaling of that of the mirroring/Danzer complex/power complex of the Tchebyshev subdivision of some simplicial complex associated to $\Gamma$ (Corollary 2.33).
We start by directly studying connection between gamma vectors and flag spheres in Section 1. This involves an overview of Boolean decompositions and gamma positivity followed by an analysis of its behavior under edge subdivisions (Proposition 1.5) and recursive properties connected to vertex decomposability (Corollary 1.7). Afterwards, we set up background information for analyzing related geometry of the simplicial complexes giving the relevant $f$-vectors in Section 2.1. A transformation that will occur frequently is taking the poset of intervals partially ordered by inclusion in Section 2.2. Putting this together, we find a connection between Tchebyshev subdivisions (repeated edge subdivisions) of “unused coloring complexes” (Definition 2.19) and interval posets of $S$ such that $\gamma(\Delta)=f(S)$ (Corollary 2.33). Finally, we use generalizations of Tchebyshev subdivisions to cell complexes and the idea behind a proof of a result of Rowlands to study connections between $f$-vectors of flag complexes coming from gamma vector constructions and $CAT(0)$ complexes.
1 Boolean decompositions of face vectors of simplicial complexes
A property that was studied in the context of possible underlying structures for gamma positivity is symmetric Boolean decompositions (see Section 4.7 on p. 84 – 85 of [27]). Another instance where Boolean decompositions appear is in possible structures of simplicial complexes whose $f$-vectors are $h$-vectors of gamma-nonnegative flag spheres.
@@ 1 Boolean decompositions of face vectors of simplicial complexes
Definition 1.2.
A simplicial complex $\Gamma$ of dimension $d-1$ has a Boolean decomposition if it an be expressed as
$\Gamma=\{F\cup G:F\in S,G\in 2^{[d-2|F|]}\}$
for some simplicial complex $S$.
We study how this decomposition behaves under transformations often used on flag simplicial complexes applied to $\Gamma$ or $\Delta$. The result below is one indication that the latter is also natural to consider.
@@
(Constantinescu–Varbaro, Theorem 4.3 on p. 99 of [10])
This motivates a generalization of the Boolean decompositions.
Definition 1.4.
A generalized Boolean decomposition of a simplicial complex $\Gamma$ of dimension $d-1$ is a decomposition of $\Gamma$ as the disjoint union of collections of faces of the form
$\{F\cup G:F\in S,G\subset 2^{[d-2|F|]}\}$
where $S$ is either a simplicial complex or the open star of one and $G$ is a subcollection of the $2^{[d-2|F|]}$ restricted by size or containing a specific element (similar to the open star).
We use this to show the following:
Proposition 1.5.
Proof.
@@
    Suppose that $a,b\in[d]$ (i.e. the Boolean part $G$ of the decomposition). Without loss of generality, we will assume that $a<b$. Note that there is an extra new vertex $v$ involved in the “Boolean” part. Then, $S$ and faces of the form $F\cup G$ with $d-2|F|\leq a-1$ or $d-2|F|\geq b+1$ (equivalently $|F|\geq\frac{d-a+1}{2}$ or $|F|\leq\frac{d-b-1}{2}$) in general are unaffected. This means that the faces $F\in S$ to consider are those with
    $\frac{d-b}{2}\leq|F|\leq\frac{d-a}{2}.$
    These are faces which involve neither $a$ nor $b$. In general, those containing only one of $a$ or $b$ are also unaffected. The ones that used to contain both $a$ and $b$ are “doubled” with this pair replaced by $a$ and $v$ or $b$ and $v$. We continue to use the same faces $F\in S$ while subsets of $[d-2|F|]$ (equivalently of $[d]$ in general) containing both $a$ and $b$ are doubled. These come from $F\in S$ such that $b\leq d-2|F|$ (equivalently $|F|\geq\frac{d-b}{2}$). While there is a decomposition into two parts, the decomposition is different being a purely Boolean decomposition. That being said, the added part has a decomposition into two parts with the first part from $F\in S$ such that $\frac{d-b}{2}\leq|F|\leq\frac{d-a}{2}$ and subsets of $[d-2|F|]$ such that $b\leq d-2|F|$ (equivalently $\frac{d-b}{2}\leq|F|$).
@@
    Now assume that $a\in V(S)$ and $b\in[d]$. The faces of $\Gamma$ affected are those using $a\in V(S)$ and $b\in[d]$. These come from faces $F\in S$ containing $a$ and subsets $G\in 2^{[d-2|F|]}$ containing $b$. In other words, we need $d-2|F|\geq b$ (equivalently $b\leq d-2|F|$ or $|F|\leq\frac{d-b}{2}$). For these faces, the subdividing vertex $v$ doubles the faces by replacing $a$ with $v$ or $b$ with $v$. The added part (faces of $\Gamma^{\prime}$ not counted by $\Gamma$) has something similar to a Boolean decomposition where we consider faces $F\in S$ that contain $a$ instead of all faces of $S$ and subsets of $[d-|F|]$ containing $b$ instead of all of $[d-2|F|]$.
@@
We now move to the other side and take $\Delta$ to be some flag Cohen–Macaulay simplicial complex. Write $h(\Delta)=f(\Gamma)$ for some (balanced) simplicial complex $\Gamma$ (which exists by work of Caviglia–Constantinescu–Varbaro [8] and Björner–Frankl–Stanley [6]). Given any flag simplicial complex $\Gamma$, there is a balanced flag vertex decomposable simplicial complex $\Delta$ such that $h(\Delta)=f(\Gamma)$ by a result of Constantinescu–Varbaro (Proposition 4.1 on p. 98 – 99 of [10]). Let $\Delta^{\prime}$ be the (stellar) edge subdivision of $\Delta$ at an edge $e\in\Delta$. Suppose that there are simplicial complexes $\Gamma$ and $\Gamma_{e}$ with Boolean decompositions
$\Gamma=\{F\cup G:F\in S,G\in 2^{[d-2|F|]}\}$
and
$\Gamma_{e}=\{F_{e}\cup G_{e}:F_{e}\in S_{e},G_{e}\in 2^{[d-2|F_{e}|-2]}\}$
involving simplicial complexes $S$ of dimension $\frac{d}{2}-1$ and $S_{e}$ of dimension $\frac{d}{2}-2$ such that $f(\Gamma)=h(\Delta)$ and $f(\Gamma_{e})=h(\operatorname{lk}_{\Delta}(e))$. This would mean that $\gamma_{i}(\Delta)=f_{i-1}(S)$ and $\gamma_{i}(\operatorname{lk}_{\Delta}(e))=f_{i-1}(S_{e})$. Note that $\gamma_{\Delta^{\prime}}(t)=\gamma_{\Delta}(t)+t\gamma_{\operatorname{lk}_{\Delta}(e)}(t)$. In other words, we have that $\gamma_{\Delta^{\prime}}(t)=f_{S}(t)+tf_{S_{e}}(t)$.
Since $\Delta$ is Cohen–Macaulay, we have that $h_{i}(\operatorname{lk}_{\Delta}(e))\leq h_{i}(\Delta)$ (Corollary 9.2 on p. 127 of [30]). This means that $f_{i-1}(S_{e})\leq f_{i-1}(S)$ and we can treat $S_{e}$ as a subcomplex of $S$ by using compression complexes. Looking at the construction for simplicial complexes whose $f$-vectors yield $h$-vectors of flag vertex decomposable simplicial complexes from Theorem 3.3 on p. 96 – 97 of [10], we see that this implies the existence of a Boolean decomposition for a simplicial complex with the same $f$-vector as $\Gamma^{\prime}$ by setting $S^{\prime}=S\cup(S_{e}*u)$ for some disjoint new vertex $u$. Going backwards from $\Gamma^{\prime}$ to $\Gamma$ is the same as removing $u$ (equivalently identifying $u$ with some vertex of $\Gamma_{e}$).
In summary, we have the following:
Proposition 1.6.
Let $\Gamma_{1}$ be a simplicial complex of dimension $D-1$ and $\Gamma_{2}\leq\Gamma_{1}$ be subcomplex of dimension $D-2$. Let $u$ be a new disjoint vertex. Note that we can instead take $\Gamma_{2}$ to be a simplicial complex such that $f_{i}(\Gamma_{2})\leq f_{i}(\Gamma_{1})$ for all $i$ since we can take compression complexes. If $\Gamma_{1}$ and $\Gamma_{2}$ have Boolean decompositions, the simplicial complex
$\Gamma\coloneq\Gamma_{1}*_{\Gamma_{2}}u=\Gamma_{1}\cup(\Gamma_{2}*u)$
also has a Boolean decomposition.
Proof.
Let
$\Gamma_{1}=\{F_{1}\cup G_{1}:F_{1}\in S_{1},G_{1}\in 2^{[d-2|F_{1}|]}\}$
and
$\Gamma_{2}=\{F_{2}\cup G_{2}:F_{2}\in S_{2},G_{2}\in 2^{[d-2|F_{2}|-1]}\}$
$\Gamma_{2}*u=\{F_{2}\cup\widetilde{G}_{2}:F_{2}\in S_{2},\widetilde{G}_{2}\in 2^{[d-2|F_{2}|]}\}.$
This yields the Boolean decomposition
$\Gamma=\{F\cup G:F_{1}\in S_{1}\cup S_{2},G\in 2^{[d-2|F|]}\}$
for $\Gamma$.
We can use this result and to apply the reasoning used in the edge subdivision case to geometrically natural examples. The statements below follow from the reasoning above and the proofs of Theorem 3.3 on p. 96 – 97 and Theorem 4.3 on p. 99 – 100 of [10].
Corollary 1.7.
  1. 1.
    Suppose that a Cohen–Macaulay simplicial complex $\Delta$ has an associated simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition. Then, the same is true for an edge subdivision $\Delta^{\prime}$ of $\Delta$. In other words, there is a simplicial complex $\Gamma^{\prime}$ such that $h(\Delta^{\prime})=f(\Gamma^{\prime})$ and $\Gamma^{\prime}$ has a Boolean decomposition.
  2. 2.
@@
    Any simplicial complexes associated to antistars and links of vertices with Boolean decompositions can be used to construct a flag (resp. quasi-flag from Definition 4.2 on p. 99 of [10]) simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition.
2 Geometry via Tchebyshev transformations and subdivisions
Using the expression of the gamma polynomial as an inverted Chebyshev expansion and work of Hetyei [18] on Tchebyshev subdivisions, we can give geometric information on simplicial complexes $T$ such that $\gamma(\Delta)=f(T)$.
2.1 Background related to cell and cubical complexes

### v2

@@ Abstract
One of the most common and effective methods of obtaining structural information on simplicial complexes is to use tools from algebraic geometry/commutative algebra (often motivated by properties of toric varieties). However, there is no general algebro-geometric description of components of the gamma vector holding for arbitary flag simplicial spheres. This invariant occurs in many different contexts including permutation statistics, signatures of toric varieties, and Euler characteristics of nonpositively curved piecewise Euclidean manifolds. Combinatorial methods resulting from an explicit inverted Chebyshev expansion give rise to new positivity properties and cell complex structures that are of interest in their own right. Note that the focus is on the $f$-vector rather than the $h$-vector in “algebraic” settings. For flag simplicial spheres $\Delta$, the fact that $h(\Delta)=f(\Gamma)$ and compatibility between Chebyshev expansions and a modification of the $f$-polynomial by work of Hetyei are the key inputs. In the main formula implying new positivity results, local structures of $CAT(0)$ complexes and cubical analogues of barycentric subdivisions give deeper connections with cubical complex structures complementing earlier work related to the top gamma vector component. Afterwards, we return to the motivating example of barycentric subdivisions and consider how $f$-vectors of Cohen–Macaulay and vertex decomposable flag complexes in geometric settings decompose and interact with geometric transformations. This includes subdivisions of simplicial complexes and recursive properties they share with vertex decomposable flag complexes.
@@ Introduction
In general, structural information on faces of simplicial complexes is often obtained via tools from commutative algebra and algebraic geometry to extract algebraic/geometric information on simplicial complex structures. This makes use of the $h$-vector (analogous to the even degree Betti numbers of simplicial toric varieties) instead of the $f$-vector that literally counts the number of faces of each dimension. Note that the two are technically related by an invertible linear transformation (p. 213 of [7]). This has been useful in many settings such as the proof of the upper bound theorem (Corollary 5.3 on p. 141 of [36], Corollary 3.5 on p. 60 of [33], Corollary 5.4.7 on p. 240 of [7], survey in [14]). However, our main object of interest is a vector constructed out of the $h$-vector where a suitable description only exists for one component.
To be more specific, this is the gamma vector of a reciprocal/palindromic polynomial. It appears in a wide range of combinatorial and geometric contexts including permutation statistics, signatures of toric varieties, and Euler characteristics of nonpositively curved piecewise Euclidean manifolds [3]. The nonnegativity of the gamma vector lies somewhere between unimodality and real-rootedness. Many of these examples fall under the setting of flag simplicial spheres. These are simplicial spheres where the 1-skeleton is the clique complex of a graph (equivalently having minimal nonfaces of size 2, see p. 100 of [33]). The top component is the only component where there is a large class of cases with an algebraic interpretation (as the signature of a toric variety by work of Leung–Reiner [23]). We also note that the top component is the source of connections with nonpositively curved piecewise Euclidean manifolds by work of Charney–Davis [9].
As it turns out, a combinatorial view towards studying the gamma vector as a whole gives novel structural information of its own right. This involves both the cell complex structure and new positivity properties. Note that the $f$-vector plays a larger role here than the $h$-vector in the context of the geometric methods mentioned at the beginning. Our approach started in [29] with looking at explicit formulas via gamma vector-like invariants for “usual” basis $x^{k}+x^{d-k}$ of palindromic/reciprocal polynomials of degree $\leq d$. This led to an interpretation of the reciprocal polynomial of the gamma polynomial as a sort of inverted Chebyshev expansion (Theorem 0.1).
@@
In addition to the properties studied in [29], the inverted Chebyshev expansion perspective also yields information on cell complex structures. This comes from the fact that $h(\Delta)=f(\Gamma)$ for an auxiliary balanced simplicial complex $\Gamma$ when $\Delta$ is a flag simplicial sphere [8] and the compatibility between Chebyshev polynomials and a variant of the $f$-polynomial by work of Hetyei [19] ($F$-polynomial from Definition 1.1, Proposition 1.2).
@@
The main geometric transformation here is repeated (stellar) subdivision of all the edges in some order (Tchebyshev subdivision from Part 2 of Definition 1.1) and a formal generalization of this (Definition 1.29, Remark 1.30). This yields a sort of type A to type B transition (e.g. Coxeter complexes from type A to type B, dually permutohedra of these types from [20]). Heuristically, this is connected to transitions between “Boolean structures” (e.g. face posets of simplicial complexes) and cubical complexes (see Theorem 1.18, Proposition 1.15, Remark 1.10).
Combining this with the earlier inverted Chebyshev expansion, we show that the reciprocal polynomial of the gamma polynomial, the $f$-polynomial of the Tchebyshev subdivision $T(\mathcal{D}(\Gamma))$ of an auxiliary complex $\mathcal{D}(\Gamma)$ constructed out of a balanced simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ for our flag simplicial sphere $\Delta$, and a cubical complex (mirroring/Danzer/power complex from Defintion 1.5) with local structure given by $T(\mathcal{D}(\Gamma))$ (Corollary 1.33).
@@
(Corollary 1.33)
Recall that
We have that
where the last equality follows from Proposition 1.27.
This “local structure” construction is motivated by a concrete construction using flag simplicial spheres $\Delta$ to build $CAT(0)$-complexes with repeated local structure given by $\Delta$ (Proposition 1.2.2 on p. 11 and Proposition 1.2.3 on p. 13 of [13], p. 21 of [12]). Note that an intermediate step between transformations between $F$-polynomials and $f$-polynomials are connected to face count conversions between these complexes (p. 302 of [5]). The complex $\mathcal{D}(\Gamma)$ determining the local structure comes from unused colors of faces in “signed proper $d$-colorings” of the 1-skeleton of $\Gamma$ (“signed unused color complex” from Corollary 1.33 and Definition 1.22). In particular, the main formula involved (Corollary 1.33) implies that reciprocal polynomial of the gamma polynomial with the variable multiplied by 2 has nonnegative coefficients (Part 1 of Corollary 1.35).
Apart from Tchebyshev subdivisions and conversions to cubical complexes, another connection to cubical complex structures is a cubical analogue of barycentric subdivisions (Definition 1.12). The $f$-vector transformation taking a simplicial complex to its cubical barycentric subdivison comes from translations, which naturally occur the positivity properties mentioned above. They also appear as an intermediate step in transformations between $F$-polynomials and $f$-polynomials. We note that our motivation for considering cubical complex structures came from studying the behavior of the expression when $\Delta$ is the barycentric subdivision of a simplicial sphere, in which case we have $\gamma(\Delta)=f(T)$ for some simplicial complex $T$ by work of Nevo–Petersen–Tenner (Theorem 1.1 on p. 1365 and Corollary 5.6 on p. 1376 of [27]).
Putting this together with the framework described in the previous paragraph, we find that there is a close connection between the $f$-polynomials of the cubical complexes above and cubical barycentric subdivisions and those of “dual complexes” of $T$ (Part 2 of Corollary 1.35). The following gives new positivity properties and formal structural results (where “formal” can be dropped in cases such as barycentric subdivisions of simplicial spheres).
Corollary 0.4.
  1. 1.
    Suppose that $h(\Delta)$ is the $h$-vector of a flag sphere of odd dimension d1d-1 (i.e. $d$ is even). If $d\geq 4$, then the polynomial $g_{\Delta}(2(u+1))=(2(u+2))^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2(u+2)}\right)$ has nonnegative coefficients. After multiplying by a power of 2, we have the $f$-polynomial of the mirroring/Danzer/power complex $M(T(\mathcal{D}(\Gamma)))$ with a nonnegative constant added to it.
  2. 2.
    By the $f_{k}^{\Box}(S)$ formula below Definition 1.12, substituting the variable $w$ by $u+1$ sends the $f$-polynomial of a simplicial complex to that of its cubical barycentric subdivision. Consider the $f$-polynomial of the mirroring/Danzer/power complex $M(T(\mathcal{D}(\Gamma)))$.
    Corollary 1.33 implies that its $f$-polynomial behaves “formally” like a double cubical barycentric subdivision associated to some integer tuple. If $\gamma(\Delta)=f(T)$, we can remove the word “formal” and take the $f$-polynomial associated to a cell complex after a rescaling as the input.
Recalling that simplicial complexes $\Gamma$ such that $h(\Delta)=f(\Gamma)$ have been a key input in our steps, this is followed by a discussion of the recursive structures of such $\Gamma$ in Section 2. In the barycentric subdivision case that led us to study connections between gamma vectors and $f$-vectors, the following a main underlying structure used.
@@
This includes tracking changes under edge subdivisions and studying the vertex decomposable case.
@@
(Proposition 2.5, Corollary 2.7)
@@
    (Proposition 2.5) The edge subdivision of a simplicial complex that has a Boolean decomposition has a generalized Boolean decomposition (Definition 2.2).
@@
    (Part 1 of Corollary 2.7) Suppose that a simplicial complex $\Delta$ has an associated simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition. Let $\Delta^{\prime}$ be the (stellar) subdivision of $\Delta$ with respect to an edge $e\in\Delta$. Then, there is a simplicial complex $\Gamma^{\prime}$ such that $h(\Delta^{\prime})=f(\Gamma^{\prime})$ and $\Gamma$ has a generalized Boolean decomposition. If $\operatorname{lk}_{\Delta}(e)$ also has $h(\operatorname{lk}_{\Delta}(e))=f(\Gamma_{e})$ for a simplicial complex $\Gamma_{e}$ with a Boolean decomposition, then $\Gamma*_{\Gamma_{e}*v}u$ and $\operatorname{ast}_{\Gamma}(p)*_{\Gamma_{e}}u$ also have Boolean decompositions for any “Boolean” vertices $u$, $v$, and $p$ of $\Gamma$.
@@
    Then, these properties are also satisfied by $\operatorname{ast}_{\Delta}(v)$ and $\operatorname{lk}_{\Delta}(v)$ in the respective cases. Also, the existence of Boolean decompositions for simplicial complexes $\Gamma_{1}$ and $\Gamma_{2}$ such that $h(\operatorname{ast}_{\Delta}(v))=f(\Gamma_{1})$ and $h(\operatorname{lk}_{\Delta}(v))=f(\Gamma_{2})$ would imply that one exists for $\Gamma\coloneq\Gamma_{1}*_{\Gamma_{2}}u$ for a new vertex $u$ (noting that $h(\Delta)=f(\Gamma)$).
Recursive properties common to both settings are listed in Proposition 2.6).
We start our approach to studying connections between gamma vector and face structures with a review of Tchebyshev transforms/subdivisions and compatibility relations in Section 1.1. This is followed by setting up background information for analyzing related geometry of the simplicial complexes giving the relevant $f$-vectors in Section 1.2.1. A transformation that will occur frequently is taking the poset of intervals partially ordered by inclusion in Section 1.2.2. Putting this together, we find a connection between Tchebyshev subdivisions (repeated edge subdivisions) of “unused coloring complexes” (Definition 1.22) and interval posets of $S$ such that $\gamma(\Delta)=f(S)$ (Corollary 1.35). Finally, we use generalizations of Tchebyshev subdivisions to cell complexes and the idea behind a proof of a result of Rowlands to study connections between $f$-vectors of flag complexes coming from gamma vector constructions and $CAT(0)$ complexes. Afterwards, we give an overview of Boolean decompositions and gamma positivity followed by an analysis of its behavior under edge subdivisions (Proposition 2.5) and recursive properties connected to vertex decomposability (Corollary 2.7).
1 Geometry of the gamma vector as a whole via Tchebyshev transformations and subdivisions
While $h$-vectors are typically perceived as more “geometric” adapting information from toric varieties, we find that the combinatorial methods of thinking about $f$-vectors and the inverted Chebyshev expansion from Theorem 0.1 gives interesting structural information from a different perspective. In particular, we obtain new positivity properties and underlying cell complex structure describing the gamma vector as a whole. The main formula obtained is in Corollary 1.33 and implications including positivity and structural properties (e.g. repeated edge subdivisions, cubical analogues of barycentric subdivisions) are given in Corollary 1.35. This is enabled by the compatibility of $f$-vector data with repeated edge subdivisions encoded by linear transformations induced by Chebyshev polynomials studied in earlier work of Hetyei. We also have a deeper connection with cubical complexes beyond the top component via transformations from faces of flag simplicial spheres $\Delta$ to $CAT(0)$-complexes with repeated local structure given by $\Delta$ (Remark 1.7). Along the way, we find a sort of conversion from type A to type B data (keeping Coxeter complexes in mind) that seems analogous to switching between face posets of simplicial and cubical complexes (e.g. multiplication by 2 involved). Following initial definitions from Section 1.1 and Section 1.2.1, the case of barycentric subdivisions of flag spheres is used to introduce and motivate transitions between “Boolean structures” (e.g. face poset of a simplicial complex) and cubical complexes and cubical barycentric subdivisions involved in Section 1.2.2. We note that the relationship between $f$-polynomials and $F$-polynomials is connected to both the local structure discussion above and cubical barycentric subdivisions giving cubical complexes.
1.1 Background on Tchebyshev transforms/subdivisions and compatibility
Before we study structural properties of cell complexes connected to Tchebyshev transformations, we first review some key definitions and properties mentioned in the introduction.
Definition 1.1.
(Hetyei, Definition 1.1 on p. 571 and Definition 2.1 on p. 574 of [19], Definition 2 on p. 921 of [20])
  1. 1.
    Given a polynomial $F(x)=a_{n}x^{n}+\ldots+a_{1}x+a_{0}\in\mathbb{R}[x]$, the Tchebyshev transforms $T(F)(x)$ and $U(F)(x)$ of the first and second kind of $F(x)$ are given by
    $T(F)(x)\coloneq a_{n}T_{n}(x)+\ldots+a_{1}T_{1}(x)+a_{0}$
    and
    $U(F)(x)\coloneq a_{n}U_{n-1}(x)+\ldots+a_{1}U_{0}(x),$
    where $T_{m}(x)$ is Tchebyshev polynomial of the first kind, determined by
    $T_{m}(\cos(\alpha))=\cos(m\cdot\alpha)$
    and $U_{m}(x)$ is the Tchebyshev polynomial of the second kind for all $m\geq 0$.
  2. 2.
    We define a Tchebyshev triangulation $T(\Delta)$ of a simplicial complex $\Delta$ having $m=f_{1}(\Delta)$ as follows. Number the edges $e_{1},\ldots,e_{m}$ in some order. We subdivide the edge $e_{i}$ into a path of length 2 by adding the midpoint $w_{i}$ and we also subdivide all the faces containing $e_{i}$ by performing a stellar subdivision. Note that the $f$-vector from any initial ordering yields the same $f$-vector (and thus the same $F$-vector) (Theorem 3 on p. 922 of [20]).
The Tchebyshev subdivision is compatible with a modification of the $f$-polynomial of a simplicial complex.
Proposition 1.2.
(Hetyei, Hetyei–Nevo from Proposition 3.3 on p. 578 – 579 and Proposition 4.4 on p. 580 – 581 of [19], Proposition 5.1 on p. 99 of [21])
Given a simplicial complex $S$, let $F_{S}(x)\coloneq f_{S}\left(\frac{x-1}{2}\right)$. Here, $f_{\Delta}(t)$ is the usual $f$-polynomial of a simplicial complex $\Delta$.
  1. 1.
    The Tchebyshev transform of the $F$-polynomial of a simplicial complex is the $F$-polynomial of its Tchebyshev triangulation:
    $T(F_{\Delta})(x)=F_{T(\Delta)}(x).$
  2. 2.
    The Tchebyshev transform of the second kind of the $F$-polynomial of a simplicial complex is half of the $F$-polynomial of its Tchebyshev triangulation:
    $U(F_{\Delta})(x)=\frac{1}{2}F_{U(\Delta)}(x).$
    Note that a Tchebyshev subivision is also the same transformation taking the type $A$ Coxeter complex to the type $B$ Coxeter complex [20]. Dually, one can consider faces of type A and type B permutohedra (p. 919 of [20]).
1.2 Cell complex structures induced by the $f$-vector perspective
1.2.1 Background related to cell and cubical complexes